# Evaluating the World Model Implicit in a Generative Model

### v2

@@ 5 Conclusion
In order to build high-fidelity algorithms that meaningfully capture the logic of the problems they model, we need ways to measure how close we are to that goal. This paper suggests theoretically grounded metrics for assessing the world models implicit inside generative models. Applications to maps, games, and logic puzzles suggest these metrics are both feasible to implement and insightful. Our primary limitation is the focus on deterministic finite automata. While it is suitable for many applications like games, logic, and state tracking, extending it would be quite valuable, e.g. to situations where the underlying world model is more complicated than a DFA or is unknown. We suspect that the core ideas related to sequence compression and sequence distinction generalize to these richer settings, but leave that to future work.
@@ Acknowledgements
Keyon Vafa is supported by the Harvard Data Science Initiative. Justin Chen is supported by an NSF Graduate Research Fellowship under Grant No. 174530. Jon Kleinberg is supported in part by a Vannevar Bush Faculty Fellowship, a Simons Collaboration grant, and a grant from the MacArthur Foundation. We thank the Chicago Booth School of Business for generous support. We thank Foundry [^4] https://www.mlfoundry.com/ for providing the compute required to conduct this research. We also thank Juan Carlos Perdomo and Neekon Vafa for helpful comments and feedback.
Appendix A Proofs
@@ Appendix B Reconstructed Maps
B.1 Algorithm
Algorithm 1 Graph Reconstruction from Sequences
[…]
Output: Reconstructed edges $\hat{D}$.
1:  Initialize $\hat{D}(u,v)\leftarrow\square$ for all $u,v\in V$
2:  for $i\in[n]$ do
3:     $j\leftarrow 2$ is an index into the list of directions
4:     $u\leftarrow s_{i}[0]$ is the current node, starting at the source
5:     while $s_{i}[j]\neq\texttt{end}$ do
6:        if $\nexists v\in V$ s.t. $\hat{D}(u,v)=s_{i}[j]$ then
7:           if |$|\{v\in V:\hat{D}(u,v)\neq\square\}|\geq deg$ then
9:           else if $\exists v\in V$ s.t. $D(u,v)=s_{i}[j]$ then
10:              $\hat{D}(u,v)\leftarrow s_{i}[j]$ // Add an edge from the real graph
11:           else
12:              $N(u)\leftarrow\{v\in V:u,v$ within Euclidean distance $dist\}$
15:        $j\leftarrow j+1$
16:     if $u\neq s_{i}[1]$ then
17:        FAIL to construct sequence
18:  return  $\hat{D}$
B.2 Maps
We include reconstructed maps built from $6400$ transformer-generated sequences for transformers trained on shortest paths, noisy shortest paths (modeling traffic), and random walks in Manhattan. Each sequence is generated by randomly sampling an (origin, destination) pair and then sampling the model’s traversal for each pair. Because the transformer never sees sequences of length more than 100 during training, we only sample pairs for which there exists a valid traversal in less than 100 moves. This distribution of (origin, destination) pairs still varies from the distributions used to train and evaluate each model, and the percent of traversals that are valid falls from >95% to 65-80%. We vary the constrained maximum degree between $4$ and $8$ (the true map has maximum degree $4$ and there are $8$ possible cardinal directions), and the maximum edge distance between $\nicefrac{{1}}{{2}}$ and $1$ mile.
Appendix C Deterministic Finite Automata and Myhill-Nerode
  1. 1.
    $Q$ is a finite set of states,
  2. 2.
    $\Sigma$ is a finite set of characters,
  3. 3.
  4. 4.
    $q_{0}\in Q$ is the start state,
  5. 5.
    $F\subseteq Q$ is the set of accepting states.
Definition C.1 (Equivalent sequences).
Theorem C.2 ([22, 23]).
Appendix D Evaluation metric details and ablations
D.1 Implementation details
For the distinction test, we sample two distinct states, $q_{1}$ and $q_{2}$, uniformly at random, and sample sequences that lead to each state, $s_{1}$ and $s_{2}$, as before. As before, computing full Myhill-Nerode boundaries is intractable, so we approximate them. To approximate the true world model’s Myhill-Nerode boundary, we consider all continuations of length $k$, and find the set of minimal suffixes that are accepted after $q_{1}$ but not $q_{2}$. To see which elements in the true boundary are distinguished by the model, we evaluate the model on each element in conditioned on $s_{1}$ vs $s_{2}$. We use $k=5$, and below we show that results are robust across different $k$. We again approximate the transformer’s Myhill-Nerode boundary by taking $M=30$ Monte-Carlo samples. For each sequence that’s accepted after $s_{1}$ but not $s_{2}$, we find the minimal distinguishing suffix and include it in the boundary set. We then assess precision by calculating which elements in the model’s boundary are acceptable after $q_{1}$ but not $q_{2}$
D.2 Ablations
We begin by considering $\epsilon$. $\epsilon$ dictates a tradeoff between precision and recall. Theoretically, larger values of $\epsilon$ improve precision and hurt recall, while smaller values hurt precision but improve recall. In the main result, we consider $\epsilon=0.01$. Table 3 reports results for other values of $\epsilon$ on the maps metrics. Empirically we see the tradeoff between precision and recall as $\epsilon$ changes. However, the conclusions are stable: every model has an incoherent world model across values of $\epsilon$. For example, while the random walks model has high compression and distinction precision for $\epsilon=10^{-6}$, it has a very low distinction recall, of $0.13$. Meanwhile, while the distinction recall is bumped up to $0.96$ for $\epsilon=0.10$, its compression precision falls to $0.44$.
Compression precision Distinction precision Distinction recall
Shortest paths 0.11 (0.03) 0.38 (0.05) 0.21 (0.03)
$\epsilon=0.10$ Noisy shortest paths 0.08 (0.03) 0.37 (0.05) 0.20 (0.03)
Random walks 0.44 (0.05) 0.99 (0.01) 0.96 (0.02)
Shortest paths 0.19 (0.01) 0.36 (0.01) 0.26 (0.01)
$\epsilon=10^{-2}$ Noisy shortest paths 0.07 (0.01) 0.36 (0.01) 0.25 (0.01)
Random walks 0.68 (0.02) 0.99 (0.00) 1.00 (0.01)
Shortest paths 0.66 (0.05) 0.36 (0.05) 0.36 (0.04)
$\epsilon=10^{-4}$ Noisy shortest paths 0.20 (0.04) 0.28 (0.04) 0.31 (0.04)
Random walks 0.86 (0.04) 1.00 (0.01) 0.99 (0.01)
Shortest paths 0.94 (0.02) 0.55 (0.05) 0.20 (0.03)
$\epsilon=10^{-6}$ Noisy shortest paths 0.67 (0.05) 0.49 (0.05) 0.17 (0.03)
Random walks 1.00 (0.00) 1.00 (0.00) 0.13 (0.03)
Our metrics approximates the true Myhill-Nerode boundary by truncating the length of suffixes it enumerates. Only the recall metric is affected by the suffix length. Figure 23 shows the recall metric for the shortest paths model for different values of the suffix length. There is a steep difference between suffix lengths of 1 and 3. However, there does not appear to be much difference past 3. This may suggest that the true Myhill-Nerode boundary for most states don’t include many sequences longer than 5. We therefore use 5 as the maximum suffix length considered for our approximation of the boundary.
Figure 23: Distinction recall performance as a function of the maximum suffix length used to approximate the true Myhill-Nerode boundary. Results are for the shortest paths map model.
Our test also relies on Monte-Carlo sampling the model’s Myhill-Nerode boundary. The number of samples affects only the precision test. A prefix pair fails the compression test whenever there’s one suffix that the model accepts for one prefix and not for the other, so performance should worsen as the number of samples increases. Table 4 shows how the precision scores vary as a function of the number of samples for the shortest paths model. As we increase the number of samples from 10 to 20, the compression precision score worsens; however, it seems to stabilize afterwards. On the other hand, there does not seem to be much of an effect on the precisino for distinction. We use $M=30$ Monte Carlo samples for our tests.
Number of samples Compression precision Distinction precision
10 0.37 (0.05) 0.40 (0.05)
20 0.17 (0.04) 0.41 (0.05)
30 0.19 (0.01) 0.36 (0.01)
Appendix E Rides data construction
Here we describe the rides dataset construction in more detail. The original dataset consists of 12,615,757 rides that took place in Manhattan between January and March 2014. After dropping duplicate rides, we’re left with 3,358,737 sequences. We use the OSMnx library [5] to represent New York as a weighted graph. We match pickups and dropoffs to the closest intersection, measured in terms of latitude/longitude. The graph consists of 4,580 nodes, 9,846 edges, and each node has a median of 2 valid intersections. We convert bearings to one of 8 cardinal directions. We remove short traversals with two node or less. We also remove sequences with more than 100 tokens.
Shortest paths.  The first approach creates traversals between two nodes by finding the shortest path between them. For each taxi ride in the dataset, we map the pickup latitude/longitude to the closest intersection and do the same for the dropoff location. We then perform Dijkstra’s algorithm to find the shortest path weighted by distance. After filtering out duplicated traversals, we have a training set of 2,932,675 sequences and 120,400,201 tokens, along with a validation set 1,000 sequences and 41,641 tokens.
Appendix F Additional results
Table 5 includes additional analysis about the capabilities of the model. We report the fraction of held-out (origin, destination) pairs for which the shortest paths and noisy shortest paths recover a near-shortest path. For the noisy shortest paths model, we report the percent that find the shortest path over all graphs used for sampling. We find that both models are capable at finding near-shortest paths.
Within 1% of shortest path Within 10% of shortest path
Shortest paths 0.98 (0.00) 0.98 (0.00)
Noisy shortest paths 0.62 (0.02) 0.88 (0.01)
Table 5: The fraction of held-out (origin, destination) pairs for which each model recovers a near-shortest path (higher is better). Standard errors are in parentheses.
Table 6 shows the results of our tests on models trained to play Othello. We use the model checkpoints provided by Li et al. [17]. We perform 1000 samples of each test.
Compression precision Distinction precision Distinction recall
Untrained transformer 0.00 (0.00) 0.02 (0.00) 0.14 (0.01)
Champion Othello 0.00 (0.00) 0.65 (0.01) 0.27 (0.01)
Synthetic Othello 0.98 (0.00) 0.99 (0.00) 1.00 (0.00)
True world model 1.00 1.00 1.00
Table 6: Our world model test applied to Othello sequence models. The champion and synthetic models refer to models trained on real-world tournament games and synthetic games, respectively.
In Section 4, we showed that while LLMs perform well at a logic task, they do so without having a coherent world model; their compression and distinction abilities are poor. The result in Section 4 was for $n=3$ chairs. Table 7 recreates the results for the setting with five chairs. While the overall task accuracy is slightly worse than it was for 3 chairs, the world model test results are similarly bad. Again, all models fail to compress and distinguish states.
Capabilities Proposed tests
Task accuracy Compression precision Distinction recall
Llama-3 (8b) 0.767 (0.023) 0.140 (0.064) 0.230 (0044)
Llama-3 (70b) 0.877 (0.014) 0.100 (0.095) 0.340 (0.085)
GPT-3.5 (turbo) 0.672 (0.029) 0.200 (0.126) 0.160 (0.055)
GPT-4 0.870 (0.015) 0.200 (0.126) 0.380 (0.082)
Mixtral-8x22B Instruct 0.680 (0.028) 0.100 (0.095) 0.293 (0.070)
Table 7: LLM logic puzzle performance when there are five chairs.

### v3

@@ 5 Conclusion
In order to build high-fidelity algorithms that meaningfully capture the logic of the problems they model, we need ways to measure how close we are to that goal. This paper suggests theoretically grounded metrics for assessing the world models implicit inside generative models. Applications to maps, games, and logic puzzles suggest these metrics are both feasible to implement and insightful. Our results show that generative models can perform impressive tasks with incoherent world models (e.g. provide directions for taxi rides). But this incoherence makes them fragile for other tasks (e.g. providing directions when there are detours). This incoherence is also problematic when we hope to use a generative model to learn something latent about the world in scientific domains.
Our primary limitation is the focus on DFAs. While it is suitable for many applications like games, logic, and state tracking, extending it would be quite valuable, e.g. to situations where the underlying world model is more complicated than a DFA or is unknown. We suspect that the core ideas related to sequence compression and sequence distinction generalize to these richer settings, but leave that to future work.
@@ Acknowledgements
Keyon Vafa is supported by the Harvard Data Science Initiative. Justin Chen is supported by an NSF Graduate Research Fellowship under Grant No. 174530. Jon Kleinberg is supported in part by a Vannevar Bush Faculty Fellowship, a Simons Collaboration grant, and a grant from the MacArthur Foundation. We thank the Chicago Booth School of Business for generous support. We thank Foundry [^4] https://www.mlfoundry.com/ for providing the compute required to conduct this research. We also thank Sarah Bentley, Juan Carlos Perdomo, and Neekon Vafa for helpful comments and feedback.
Appendix A Proof
@@ Appendix B Reconstructed Maps
Figure 6: Performance metrics as a function of the maximum suffix length used to approximate the Myhill-Nerode boundary.
@@
B.1 Algorithm
Algorithm 1 Graph Reconstruction from Sequences
[…]
Output: Reconstructed edges $\hat{D}$.
1:  Initialize $\hat{D}(u,v)\leftarrow\square$ for all $u,v\in V$
2:  for $i\in[n]$ do
3:     $j\leftarrow 2$ is an index into the list of directions
4:     $u\leftarrow s_{i}[0]$ is the current node, starting at the source
5:     while $s_{i}[j]\neq\texttt{end}$ do
6:        if $\nexists v\in V$ s.t. $\hat{D}(u,v)=s_{i}[j]$ then
7:           if |$|\{v\in V:\hat{D}(u,v)\neq\square\}|\geq deg$ then
9:           else if $\exists v\in V$ s.t. $D(u,v)=s_{i}[j]$ then
10:              $\hat{D}(u,v)\leftarrow s_{i}[j]$ // Add an edge from the real graph
11:           else
12:              $N(u)\leftarrow\{v\in V:u,v$ within Euclidean distance $dist\}$
15:        $j\leftarrow j+1$
16:     if $u\neq s_{i}[1]$ then
17:        FAIL to construct sequence
18:  return  $\hat{D}$
B.2 Maps
We include reconstructed maps built from $6400$ transformer-generated sequences for transformers trained on shortest paths, noisy shortest paths (modeling traffic), and random walks in Manhattan. Each sequence is generated by randomly sampling an (origin, destination) pair and then sampling the model’s traversal for each pair. Because the transformer never sees sequences of length more than 100 during training, we only sample pairs for which there exists a valid traversal in less than 100 moves. We note that this distribution of (origin, destination) pairs varies from the distributions used to train and evaluate each model, and the percent of traversals that are valid falls from >95% to 65-80%; however, Figure 12 shows similar maps when each map is constructed by sampling (origin, destination) pairs from the training/evaluation distributions. We vary the constrained maximum degree between $4$ and $8$ (the true map has maximum degree $4$ and there are $8$ possible cardinal directions), and the maximum edge distance between $\nicefrac{{1}}{{2}}$ and $1$ mile.
Appendix C Deterministic Finite Automata and Myhill-Nerode
  1. 1.
    $Q$ is a finite set of states,
  2. 2.
    $\Sigma$ is a finite set of characters,
  3. 3.
  4. 4.
    $q_{0}\in Q$ is the start state,
  5. 5.
    $F\subseteq Q$ is the set of accepting states.
Definition C.1 (Equivalent sequences).
Theorem C.2 ([26, 27]).
Appendix D Additional results
Figure 6 shows how our performance metrics vary as we consider different suffix lengths for approximating the Myhill-Nerode boundary. For compression precision, considering a boundary of size $k$ corresponds to sampling suffixes of length-$k$ for each suffix and measuring whether prefixes with the same state have the same length-$k$ suffixes. For distinction precision, we consider a boundary of size $k$ by only sampling $k$-length suffixes to approximate a model’s Myhill-Nerode boundary. For distinction recall, we consider a boundary of size $k$ by only constructing the true Myhill-Nerode boundary based on $k$-length suffixes. The results in Figure 6 show the importance of considering larger boundaries as opposed to smaller ones (e.g. single next-tokens); for example, while the model trained on random walks scores 100% on compression precision when boundaries of length-1 are considered, its performance is 50% when the full Myhill-Nerode boundary is considered.
Table 3 reports our metrics on models trained to play Othello. We use the transformer model checkpoints provided by Li et al. [20]. We perform 1000 samples of each test. Our metrics find that while the model trained on synthetic data recovers the true world model, the model trained on championship data does not. In Table 4, we perform a detour exercise analogous to the one performed for taxi rides in Section 3, where here a model’s predicted move is replaced with another legal one. The detour results support the discernment between models found by our metrics; while the model trained on synthetic data produces near-perfect games regardless of detours, the model trained on championship data fails immediately.
Compression precision Distinction precision Distinction recall
Untrained transformer 0.00 (0.00) 0.02 (0.00) 0.14 (0.01)
Championship Othello 0.00 (0.00) 0.65 (0.01) 0.27 (0.01)
Synthetic Othello 0.98 (0.00) 0.99 (0.00) 1.00 (0.00)
True world model 1.00 1.00 1.00
Table 3: The metrics from Section 2.4 applied to Othello sequence models. The championship and synthetic models refer to models trained on real-world tournament games and synthetic games, respectively.
Probability of detour
0% 1% 10% 25% 50%
Random Championship Othello 1.00 (0.00) 0.66 (0.05) 0.05 (0.02) 0.01 (0.01) 0.01 (0.01)
detours Synthetic Othello 1.00 (0.00) 0.99 (0.01) 0.97 (0.02) 0.97 (0.02) 0.99 (0.01)
True world model 1.00 1.00 1.00 1.00 1.00
Adversarial Championship Othello 1.00 (0.00) 0.70 (0.05) 0.01 (0.01) 0.01 (0.01) 0.00 (0.00)
detours Synthetic Othello 1.00 (0.00) 0.98 (0.01) 0.99 (0.01) 0.96 (0.02) 0.97 (0.02)
True world model 1.00 1.00 1.00 1.00 1.00
Table 4: The fraction of traversals that are valid when detours are introduced in Othello. For “random detours”, a model’s proposed token is replaced with a randomly chosen (true) valid token; for “adversarial detours”, it is replaced with the model’s lowest ranked valid token.
Appendix E Evaluation metric details and ablations
E.1 Implementation details
For the distinction test, we sample two distinct states, $q_{1}$ and $q_{2}$, uniformly at random, and sample sequences that lead to each state, $s_{1}$ and $s_{2}$, as before. As before, computing full Myhill-Nerode boundaries is intractable, so we approximate them. To approximate the true world model’s Myhill-Nerode boundary, we consider all continuations of length $k$, and find the set of minimal suffixes that are accepted after $q_{1}$ but not $q_{2}$. To see which elements in the true boundary are distinguished by the model, we evaluate the model on each element in conditioned on $s_{1}$ vs $s_{2}$. We use $k=5$, and find that results are robust across different $k\geq 5$. We again approximate the transformer’s Myhill-Nerode boundary by taking $M=30$ Monte-Carlo samples. For each sequence that’s accepted after $s_{1}$ but not $s_{2}$, we find the minimal distinguishing suffix and include it in the boundary set. We then assess precision by calculating which elements in the model’s boundary are acceptable after $q_{1}$ but not $q_{2}$
E.2 Ablations
We begin by considering $\epsilon$, which dictates a tradeoff between precision and recall. In the main text, we consider $\epsilon=0.01$. Table 5 reports results for other values of $\epsilon$ on the maps metrics. Empirically we see the tradeoff between precision and recall as $\epsilon$ changes. However, the conclusions are stable: every model has an incoherent world model across values of $\epsilon$. For example, while the random walks model has high compression and distinction precision for $\epsilon=10^{-6}$, it has a very low distinction recall, of $0.11$. Meanwhile, while the distinction recall is bumped up to $1.00$ for $\epsilon=0.10$, its compression precision falls to $0.16$.
Compression precision Distinction precision Distinction recall
Shortest paths 0.08 (0.03) 0.30 (0.05) 0.16 (0.03)
$\epsilon=0.10$ Noisy shortest paths 0.04 (0.02) 0.36 (0.05) 0.20 (0.03)
Random walks 0.16 (0.04) 1.00 (0.00) 1.00 (0.00)
Shortest paths 0.10 (0.03) 0.38 (0.05) 0.18 (0.03)
$\epsilon=10^{-2}$ Noisy shortest paths 0.08 (0.03) 0.38 (0.05) 0.26 (0.03)
Random walks 0.46 (0.06) 1.00 (0.00) 1.00 (0.00)
Shortest paths 0.36 (0.05) 0.29 (0.05) 0.29 (0.04)
$\epsilon=10^{-4}$ Noisy shortest paths 0.29 (0.05) 0.35 (0.05) 0.34 (0.04)
Random walks 0.79 (0.05) 1.00 (0.00) 0.97 (0.02)
Shortest paths 0.70 (0.05) 0.29 (0.05) 0.29 (0.04)
$\epsilon=10^{-6}$ Noisy shortest paths 0.70 (0.05) 0.48 (0.05) 0.23 (0.04)
Random walks 0.99 (0.01) 0.99 (0.01) 0.11 (0.03)
The metrics introduced in Section 2.4 depend on defining what it means for a model to accept or reject a sequence. In the main text we consider acceptance based on a threshold parameter, which corresponds to an $\epsilon$-sampling decoding mechanism [12]. Here, we consider two alternative forms of acceptance based on other decoding strategies: top-$k$ [8] and top-$p$ [13]. For acceptance based on top-$k$, a token is accepted if it’s in the model’s top-$k$ ranked tokens for a sequence and rejected otherwise. For top-$p$, a token is accepted if it’s part of the the smallest set of highest-probability tokens whose cumulative probability is larger than $p$. Results are depicted in Table 6 and Table 7 and point to the same conclusion as the threshold-based metrics in Section 3; none of the models have recovered the world model, but the model trained on random walks performs best.
Compression precision Distinction precision Distinction recall
Shortest paths 0.14 (0.04) 0.33 (0.05) 0.17 (0.03)
$k=1$ Noisy shortest paths 0.06 (0.04) 0.35 (0.07) 0.11 (0.03)
Random walks 0.40 (0.08) 0.69 (0.06) 0.30 (0.04)
Shortest paths 0.21 (0.05) 0.32 (0.05) 0.17 (0.03)
$k=2$ Noisy shortest paths 0.07 (0.04) 0.31 (0.07) 0.23 (0.05)
Random walks 0.21 (0.07) 0.93 (0.03) 0.73 (0.05)
Shortest paths 0.49 (0.06) 0.41 (0.05) 0.30 (0.03)
$k=4$ Noisy shortest paths 0.44 (0.08) 0.22 (0.06) 0.33 (0.06)
Random walks 0.64 (0.08) 0.98 (0.02) 0.51 (0.06)
Table 6: Compression and distinction metrics for maps data where token acceptance is based on the top-$k$ decoding mechanism. We consider values from $k=1$ to $k=4$ because there are only 8 valid cardinal directions.
Compression precision Distinction precision Distinction recall
Shortest paths 0.05 (0.02) 0.33 (0.05) 0.17 (0.03)
$p=0.90$ Noisy shortest paths 0.08 (0.04) 0.34 (0.07) 0.20 (0.05)
Random walks 0.16 (0.06) 0.96 (0.04) 0.94 (0.03)
Shortest paths 0.22 (0.05) 0.39 (0.05) 0.21 (0.03)
$p=0.99$ Noisy shortest paths 0.03 (0.03) 0.39 (0.07) 0.24 (0.05)
Random walks 0.54 (0.08) 1.00 (0.00) 1.00 (0.00)
Shortest paths 0.31 (0.05) 0.35 (0.05) 0.22 (0.04)
$p=0.999$ Noisy shortest paths 0.15 (0.06) 0.46 (0.07) 0.31 (0.05)
Random walks 0.73 (0.07) 1.00 (0.00) 1.00 (0.00)
Table 7: Compression and distinction metrics for maps data where token acceptance is based on the top-$p$ decoding mechanism.
Our test also relies on Monte-Carlo sampling the model’s Myhill-Nerode boundary. The number of samples affects only the precision test. A prefix pair fails the compression test whenever there’s one suffix that the model accepts for one prefix and not for the other, so performance should worsen as the number of samples increases. Table 8 shows how the precision scores vary as a function of the number of samples for the shortest paths model. We use $M=30$ Monte Carlo samples for our main reported metrics.
Number of samples Compression precision Distinction precision
10 0.14 (0.04) 0.35 (0.05)
20 0.12 (0.04) 0.35 (0.05)
30 0.10 (0.03) 0.30 (0.04)
Appendix F Rides data construction and training
Here we describe the rides dataset construction in more detail. Our empirical studies are based on a dataset of taxi rides in New York from 2014, originally released by the NYC Taxi & Limousine Commission. We use a subset of 15 million rides that took place between January and March 2014, made available by Murray [25]. We drop duplicate rides and subset the dataset to only include rides in Manhattan, resulting in 3,358,737 sequences. We use the OSMnx library [5] to represent New York as a weighted graph. We match pickups and dropoffs to the closest intersection, measured in terms of latitude/longitude. The graph consists of 4,580 nodes, 9,846 edges, and each node has a median of 2 valid intersections. We convert bearings to one of 8 cardinal directions. We remove short traversals with two node or less. We also remove sequences with more than 100 tokens.
Shortest paths.  The first approach creates traversals between two nodes by finding the shortest path between them. For each taxi ride in the dataset, we map the pickup latitude/longitude to the closest intersection and do the same for the dropoff location. We then perform Dijkstra’s algorithm to find the shortest path weighted by distance. After filtering out duplicated traversals, we have a training set of 2,932,675 sequences and 120,400,201 tokens, along with a validation set of 1,000 sequences and 41,641 tokens.
We use the GPT-2 architecture [29] to train models on all datasets. We use the GPT-2 small architecture for the shortest paths model and the GPT-2 extra-large architecture for the noisy shortest paths and random walks models. For the shortest paths models, we train until we overfit and use the best validation checkpoint. For the two larger datasets, we train for a fixed number of epochs and use the last validation checkpoint; we use 5 epochs for the noisy shortest paths dataset and 1 epoch for the random walks dataset. We train all models on 8 A100 GPUs, using a batch size of 6 sequences per GPU. Training time ranges from about 12 hours for the shortest paths model to 48 hours for the random walks model.
Appendix G Additional maps
@@
(a) Shortest paths
(b) Noisy shortest paths
(c) Random walks
Figure 12: Reconstructed maps from transformers trained on shortest paths, noisy shortest paths, and random walks constructed using (origin, destination) pairs sampled from the same distribution used for training. While trajectories are valid more than 95% of the time, the errors reveal incoherence. Reconstruction uses 50k sequences with maximum degree 4 and maximum edge length $\nicefrac{{1}}{{2}}$ mile.